A chain store manager has been told by the main office that daily profit, P, is related to the number of clerks working that day, x, according to the equation P = -25x^2 + 450x. What number of clerks will maximize the profit?
Answer by ewatrrr(24785)

 
Hi,
y = ax^2 + bx + c ⇒ y = a(x -(-b/2a)^2) - a(-b/2a)^2 + c
P = -25x^2 + 450x. | Note: -450/(2*-25) = 9
P= -25(x-9)^2 + 2025
9 cleks will maximize the daily profit at $2025